# Methods and Open Problems in Differentiable Social Choice: Learning Mechanisms, Decisions, and Alignment

### v2

@@ Abstract
Social choice is no longer a peripheral concern of political theory or economics—it has become a foundational component of modern machine learning systems. From auctions and resource allocation to federated learning, participatory governance, and the alignment of large language models, machine learning pipelines increasingly aggregate heterogeneous preferences, incentives, and judgments into collective decisions. In effect, many contemporary machine learning systems already implement social choice mechanisms, often implicitly and without explicit normative scrutiny.
@@
This Review synthesizes this emerging paradigm—what we term differentiable social choice—across auctions, voting rules, participatory budgeting, liquid democracy, decentralized aggregation, AI alignment, and inverse mechanism learning. We articulate a unifying perspective in which loss functions act as implicit aggregation rules, architectural inductive biases encode axioms such as anonymity or feasibility, and properties like robustness, proportionality, and incentive compatibility emerge as learned—but fragile—outcomes. We show how classical impossibility results and axiomatic trade-offs do not disappear under learning, but are relocated into objectives, constraints, data collection protocols, and optimization dynamics.
Beyond synthesizing methods, this Review is explicitly forward-looking. Across six domains, we identify 36 concrete open problems spanning incentive guarantees, robustness to distribution shift and strategic manipulation, certification and auditability of learned mechanisms, pluralistic preference aggregation, governance of alignment objectives, and the limits of inverse inference from behavior. These open problems delineate a new research agenda at the intersection of machine learning, economics, and social choice theory.
We argue that as machine learning systems increasingly mediate collective decisions at societal scale, the success of data-driven aggregation will depend not only on empirical performance, but on our ability to specify, verify, and govern the social choice principles that learning systems inevitably implement.
@@ 1 Introduction
Modern deployments have amplified these limitations. Spectrum auctions, online advertising markets, participatory budgeting, decentralized governance platforms, and preference-based training of large language models (LLMs) all operate in decision spaces that are high-dimensional, combinatorial, non-convex, and distribution-dependent. Yet, social choice is (both implicitly and explicitly) used in many of the most important contemporary machine learning research, e.g. in automated auction mechanism design for revenue optimzation Duan et al. (2023) and Large Language Models joint creative generations Duetting et al. (2024), in stake-holder portfolio balance and participatary resource allocation Fairstein et al. (2024), in federated learning where model updates are aggregated based on a certain mechanism Kesari & Kesari (2025), and, most celebrated recently, in preference-based alignment (RLHF) of LLMs from human annotators who have inherently heterogeneous preference models Siththaranjan et al. (2024); Swamy et al. (2024). We summarize these applications of social choice in machine learning in Table 1. In such settings, analytic derivations with universal guarantees can be simultaneously too pessimistic—ruling out mechanisms that perform well on realistic data—and too optimistic—failing catastrophically under distribution shift or strategic adaptation Hart et al. (2013); Caiata et al. (2025). Nevertheless, many contemporary applied systems are already learned: allocation policies, ranking rules, and reward models are trained by gradient descent, whether or not designers acknowledge their implicit social-choice content.
@@
Concrete applications already illustrate both the promise and the stakes of this shift. In auction design, neural mechanisms have been trained to approximate revenue-optimal auctions in multi-item settings where analytic solutions are unknown, encoding incentive compatibility as regret penalties or architectural constraints Dütting et al. (2019); Ivanov et al. (2022); Hertrich et al. (2023). In participatory budgeting, learned aggregation rules operate over knapsack-feasible portfolios, optimizing proportionality and efficiency objectives at scales where classical solvers struggle, while benchmarking against principled rules such as the Method of Equal Shares (MES) Peters et al. (2021); Fairstein et al. (2024). In liquid democracy, delegation graphs are implemented as differentiable aggregation architectures, directly connecting social choice to federated learning and edge aggregation under trust constraints Kesari & Kesari (2025); Alouf-Heffetz et al. (2025). In AI alignment, preference-learning objectives used to train reward models and policies have been shown to implicitly instantiate specific voting rules—most notably Borda count—even when designers intend to remain agnostic about normative aggregation Siththaranjan et al. (2024); Dai & Fleisig (2024). More recently, inverse mechanism learning has turned these tools toward governance and accountability, aiming to reconstruct opaque incentive structures from observed strategic behavior An & Du (2026); Ahmadi et al. (2023).
These developments mark a deep conceptual shift. We structurally illustrate this conceptual shift in Figure 1. Differentiable social choice does not abandon classical theory; it repurposes it. Social choice axioms become a specification language for loss functions and architectures, and an audit vocabulary for interpreting learned systems. Anonymity and neutrality translate into permutation-equivariant neural networks; feasibility and budget constraints become differentiable solver layers; incentive compatibility is relaxed into approximate regret bounds that can be empirically evaluated and stress-tested. At the same time, the differentiable turn exposes new risks and novel research gaps: soft constraints may fail silently, expressive models may hide violations, and optimization objectives may encode unintended normative commitments.
@@
Data-driven PB aggregation Citizens / stakeholders Approvals or utility vectors Machine learned selection of portfolios balances utility and representation, analogous to PB rules Learning aggregation rules in participatory budgeting (data-driven PB) Fairstein et al. (2024)
Federated learning / decentralized aggregation Distributed clients Local model updates Weighted or consensus-based aggregation corresponds to a voting/collective influence mechanism Liquid democracy as hybrid voting model (proxy delegation) Kesari & Kesari (2025)
@@
Platform governance / incentive forensics Strategic users Behavioral adaptation traces Inferring effective rules from observed strategic behavior is analogous to inverse mechanism design IC and mechanism design for sociotechnical systems in AI alignment Zhang et al. (2024); An & Du (2026)
This Review surveys this rapidly evolving landscape with a unifying perspective. We organize the literature into six interconnected domains: (1) Differentiable Economics, which studies learning-based approximations to optimal auctions and contracts; (2) Neural Social Choice, which synthesizes, analyzes, and audits voting rules using deep learning; (3) Participatory Budgeting, where AI systems support collective resource allocation under combinatorial constraints; (4) Liquid Democracy, linking delegation and decentralization to graph-based learning; (5) AI Alignment as Social Choice, which interprets preference learning and reinforcement learning from human feedback (RLHF) as implicit voting; and (6) Inverse Mechanism Learning, which seeks to infer and audit incentive structures from behavior.
Our goal is not merely to catalog methods, but to articulate a coherent narrative: the transition from axiomatic artifacts to learned institutions. As collective decision-making becomes embedded in machine learning pipelines, the design of loss functions, architectures, and evaluation protocols becomes inseparable from social choice theory. Differentiable social choice thus emerges not only as a technical research program, but as a foundational framework for the governance, auditability, and legitimacy of algorithmic decision-making systems.

### v3

# Open Problems in Differentiable Social Choice: Learning Mechanisms, Decisions, and Alignment

@@ Abstract
Social choice has become a foundational component of modern machine learning systems. From auctions to resource allocation and the alignment of large generative models, machine learning pipelines increasingly aggregate heterogeneous preferences and incentives into collective decisions. In effect, many contemporary machine learning systems already implement social choice mechanisms, often implicitly and without explicit normative scrutiny.
@@
This Review synthesizes this emerging paradigm—what we term differentiable social choice—across auctions, decision aggregation, and AI alignment. We articulate a unifying perspective in which loss functions act as implicit aggregation rules, architectural inductive biases encode axioms such as anonymity or feasibility, and properties like robustness, proportionality, and incentive compatibility emerge as learned—but fragile—outcomes.
Beyond synthesizing methods, this Review is explicitly forward-looking. Across three domains, we identify 18 concrete open problems spanning incentive guarantees, robustness to distribution shift and strategic manipulation, certification and auditability of learned mechanisms, pluralistic preference aggregation, and the governance of alignment objectives. These open problems delineate a new research agenda at the intersection of machine learning, economics, and social choice theory.
@@ 1 Introduction
Modern deployments have amplified these limitations. Spectrum auctions, online advertising markets, participatory budgeting, decentralized governance platforms, and preference-based training of large language models (LLMs) all operate in decision spaces that are high-dimensional, combinatorial, non-convex, and distribution-dependent. Yet, social choice is (both implicitly and explicitly) used in many of the most important contemporary machine learning research, e.g. in automated auction mechanism design for revenue optimzation Duan et al. (2023), in Large Language Models joint creative generations Duetting et al. (2024), and, most celebrated recently, in preference-based alignment (RLHF) of LLMs from human annotators who have inherently heterogeneous preference models Siththaranjan et al. (2024); Swamy et al. (2024). We summarize these applications of social choice in machine learning in Table 1. In such settings, analytic derivations with universal guarantees can be simultaneously too pessimistic—ruling out mechanisms that perform well on realistic data—and too optimistic—failing catastrophically under distribution shift or strategic adaptation Hart et al. (2013); Caiata et al. (2025). Nevertheless, many contemporary applied systems are already learned: allocation policies, ranking rules, and reward models are trained by gradient descent, whether or not designers acknowledge their implicit social-choice content.
@@
Concrete applications already illustrate both the promise and the stakes of this shift. In auction design, neural mechanisms have been trained to approximate revenue-optimal auctions in multi-item settings where analytic solutions are unknown, encoding incentive compatibility as regret penalties or architectural constraints Dütting et al. (2019); Ivanov et al. (2022); Hertrich et al. (2023). In AI alignment, preference-learning objectives used to train reward models and policies have been shown to implicitly instantiate specific voting rules—most notably Borda count—even when designers intend to remain agnostic about normative aggregation Siththaranjan et al. (2024); Dai & Fleisig (2024).
These developments mark a deep conceptual shift. We structurally illustrate this conceptual shift in Figure 1. Social choice axioms become a specification language for loss functions and architectures, and an audit vocabulary for interpreting learned systems. Anonymity and neutrality translate into permutation-equivariant neural networks; feasibility and budget constraints become differentiable solver layers; incentive compatibility is relaxed into approximate regret bounds that can be empirically evaluated and stress-tested. At the same time, the differentiable turn exposes new risks and novel research gaps: soft constraints may fail silently, expressive models may hide violations, and optimization objectives may encode unintended normative commitments.
@@
This Review surveys this rapidly evolving landscape with a unifying perspective. We organize the literature into six interconnected domains: (1) Differentiable Economics, which studies learning-based approximations to optimal auctions and contracts; (2) Neural Social Choice, which synthesizes, analyzes, and audits voting rules using deep learning; and (3) AI Alignment as Social Choice, which interprets preference learning and reinforcement learning from human feedback (RLHF) as implicit voting.
As collective decision-making becomes embedded in machine learning pipelines, the design of loss functions, architectures, and evaluation protocols becomes inseparable from social choice theory. Differentiable social choice thus emerges not only as a technical research program, but as a foundational framework for the safety, interpretability, and auditability of algorithmic decision-making systems.